AOB

Speaker : All ( All )


Talk time : 12:35, Duration : 00:20

A few things need to be followed upon:

  • Catch up item from Software and Computing phone meeting
    • Priority was increased for Yuri with the idea of providing more CPU for the study on aging effect - did this work?
    • PDSF library was supposed to be tested - status?
    • PMD code pending: where are we? Summary.
       
  • Production general readiness
    • VF: for the beam energy scan, we will do NOTHING else than what is being used now - whenever Rosi will be back, will resume work on VF evaluation. Likely will go with Minuit - note that the UPC folks asked for PPVF (noCTB) during the run to be used for any stream matching "upc". TOF should alo be revisited.
    • We discussed the general readiness for the calibrations a few times and accuracy of calibrations were summarized at the t STAR analysis meeting at UCLA in the S&C overview talk. A quick sumamry of the mian point then:
      • Slide 13 & calibration readiness:  FTPC now done, TOF using 39 GeV calib for 7 & 11 GeV BES, PMD 62 & 39 GeV done and 7 GeV was "in progress", EMCs using Run 9 gains and will re-asses post-production, BeamLine not done yet (affects 200 GeV)
      • TPC ... best to see slides 19 and 20 which are copies of Gene's summary available at a past meeting Calibration readiness for Run 10, achieved resolution and incoming improvements summary
      • Additional distorions TBD
  • Bugs of the week was sent by Lidia yesterday - SL10h has been released (see release history) but we are still missing
    • #1941 - PMD reader doesn't read old run 2008 daq data; this is a backward compatibility issue and without this, we have a path forward but not a final library.

    • #1913 -  Simulation has extra dead TPC region; the problem is related to tpcAnodeHVavg table entries in DB; table was populated and run 2009 problem was solved needs final check for run 2010 data;  bug is not related to software; hopefuly will be closed soon
       
  • Incomings
    • ROOT 5.26.00
      • ROOT 5.26.00 will NOT be used for the initial production but may be considered for later productions - expected gains include IO improvement and overall speed improvements (TBC)
      • Compilation ongoing (SL5 32 bits gcc 4.32 ready, 64 bits on the way) - ETA by the end of the week, we would start testing in .DEV .
      • Compilation and new test bench will be made for gcc 4.50 - if performance gain is confirmed, this will trigger a compilation of OPTSTAR and root4star with gcc 4.50. ETA + 2 weeks
    • Reminder: cleanup of /star/u will be ongoing. Please consult You do not have access to view this node for more information. Last chance before moving along.
      • Additional earlier users flagged for backup were also mentionned in this posting and included users: vernet, vanyashi, deph, hiromi, hippolyt, bombara, fretiere, cmirinov, dietel
      • New targetted users include: pbaba, mheinz, benedos, knips, supriya, srn, qattan, bchkim, kowalik, antai, yanlv, saugyc, cjones, mercedes, bombara, rparra, ebrost, guojilin, ajay, mmoura
      • Note that users with a 1 year+ expiration will be removed from Hypernews and their Drupal account disabled.